OpenStudy (gabylovesyou):

Factor completely: 2x2 − 6x − 56 2(x − 7)(x + 4) (2x − 14)(x + 4) (x − 7)(2x + 8) Prime

OpenStudy (gabylovesyou):

@HawkCrimson

OpenStudy (johnweldon1993):

well what can be factored out of all the terms?

OpenStudy (anonymous):

let her do it by herself :) ive been helping her and i want to see how far she'll get :D

OpenStudy (johnweldon1993):

ahh okay @HawkCrimson yes lets see lol

OpenStudy (anonymous):

if u r not too comfortable with actually factoring it, you could always do the equations from the answer choices, and see which one matches the pre-factored form... first FOIL then multiply it all by 2. but yes, you should try it first yourself! :)

OpenStudy (anonymous):

thx @johnweldon1993 , so @Gabylovesyou do u know how to start this?

OpenStudy (anonymous):

but factoring it normally is more efficient :)

OpenStudy (anonymous):

very true @iheartfood , and she should do that in an exam, ik i would but in case of practice then do it the long and hard way :D lol

OpenStudy (anonymous):

yup! I totally agree :)

OpenStudy (gabylovesyou):

sorry i was out... ok ummm hold up

OpenStudy (anonymous):

lol i havnt left so ok xD

OpenStudy (gabylovesyou):

how do i do it when its an exponent?

OpenStudy (anonymous):

u remember the 4 steps i taught u?

OpenStudy (anonymous):

cmon plz tell me u wrote it down some where??

OpenStudy (gabylovesyou):

-,- idk how to do this one..

OpenStudy (anonymous):

but do u have the 4 steps i showed u?

OpenStudy (gabylovesyou):

The GCF etc.. right?

OpenStudy (anonymous):

the one which starts off 1st Step: (x ) (x )

OpenStudy (anonymous):

remember that?

OpenStudy (anonymous):

ill show u again if u dont remember

OpenStudy (anonymous):

?

OpenStudy (gabylovesyou):

;/

OpenStudy (anonymous):

lol it's ok :)

OpenStudy (anonymous):

\[2x ^{2} − 6x − 56\] (2x - 14)(x + 4)

OpenStudy (anonymous):

u there? i need u to be here for me too explain it

OpenStudy (gabylovesyou):

sorry im here..

OpenStudy (anonymous):

ok u have time on ur hands?

OpenStudy (gabylovesyou):

yes

OpenStudy (anonymous):

cool :D

OpenStudy (anonymous):

Step 1: (x +)(x +)

OpenStudy (anonymous):

actually gaby, this isnt something u learn from steps, u have to practice till u get the feel of it so im just going to explain it then u'll tell me if u understand and we'll practice with more questions and if ur like me then u'll forget it by tomorrow and we'll go at it again :P

OpenStudy (anonymous):

math's just a bi.tch like that xD

OpenStudy (anonymous):

i thought u had time on ur hands?

OpenStudy (anonymous):

lol^

OpenStudy (gabylovesyou):

i am like u .... LMAO ... i do but my OS is acting stupid

OpenStudy (anonymous):

wat is OS?

OpenStudy (gabylovesyou):

openstudy

OpenStudy (anonymous):

lol that's ok, everyone's openstudy was acting dump yesturday

OpenStudy (gabylovesyou):

lol ok..

OpenStudy (anonymous):

\[ 2x ^{2} − 6x − 56\] ok so back too ur question :P

OpenStudy (gabylovesyou):

lol .... so first ?

OpenStudy (anonymous):

(x + ) (x + )

OpenStudy (anonymous):

next we list all the factors of 56 1 2 4 7 56 23 14 8

OpenStudy (anonymous):

(2x + ) (x + )

OpenStudy (anonymous):

one of these combinations will be placed in the empty spaces in (2x + ) (x + )

OpenStudy (gabylovesyou):

ok...

OpenStudy (anonymous):

so lets choose 4 and 14 (2x + 4) (x + 14)

OpenStudy (anonymous):

from here u should be able too see (2x - 14) (x + 4)

OpenStudy (anonymous):

is ur answer

OpenStudy (anonymous):

u there?

OpenStudy (gabylovesyou):

UGH OS IS KILLING ME HERE.. lol

OpenStudy (gabylovesyou):

ohh ok ...

OpenStudy (anonymous):

let me give u an easier question and we'll work up from there ok?

OpenStudy (gabylovesyou):

ok :)

OpenStudy (anonymous):

\[x ^{2}+3x+2\]

OpenStudy (anonymous):

factorize that

OpenStudy (gabylovesyou):

i need to list all the factors of 2 ?

OpenStudy (anonymous):

yep

OpenStudy (gabylovesyou):

1 2 -1 -2

OpenStudy (gabylovesyou):

i choose 2 and -2... am i going good?

OpenStudy (anonymous):

2 and -2 isnt a factor of 2. only 1 and 2 r factors of 2

OpenStudy (gabylovesyou):

-.- ok then i pick 1 and 2 lol

OpenStudy (anonymous):

yay :D

OpenStudy (anonymous):

wat is ur final answer?

OpenStudy (gabylovesyou):

(x + 1)(x + 2) ?

OpenStudy (anonymous):

yes! congrates :D

OpenStudy (gabylovesyou):

OMG YAYAYYAYAYAYAYYA ...

OpenStudy (anonymous):

:D next question now :)

OpenStudy (anonymous):

\[x ^{2}+5x+6\]

OpenStudy (anonymous):

Factorize that :))

OpenStudy (gabylovesyou):

1 2 3 6 i pick 2 and 6 (x + 2)(x + 6) ????

OpenStudy (anonymous):

no cause 2*6 = 12

OpenStudy (gabylovesyou):

so 2 and 3

OpenStudy (gabylovesyou):

(x + 2)(x + 3)

OpenStudy (anonymous):

yes :D wats ur final answer?

OpenStudy (anonymous):

yes :)) next question! this is math u do a bunch of questions :P

OpenStudy (gabylovesyou):

Factor completely: x^2 − 2x + 5 (x + 5)(x − 1) Prime (x − 5)(x − 1) (x + 5)(x + 1) <---------- ANSWER

OpenStudy (anonymous):

no cause (x + 5)(x +1) = x^2 +6x +5

OpenStudy (gabylovesyou):

hmmmm so then C ?

OpenStudy (anonymous):

x^2 -2x +5 is that a question?

OpenStudy (anonymous):

C for wat? lol now im lost xD

OpenStudy (gabylovesyou):

LOL .......is this prime?

OpenStudy (anonymous):

forget the original question, we need practice witth easier ones :P

OpenStudy (anonymous):

u have time on ur hands correct?

OpenStudy (gabylovesyou):

YES .. LOL

OpenStudy (anonymous):

so next question?

OpenStudy (gabylovesyou):

yes

OpenStudy (anonymous):

\[x ^{2}-x-12\]

OpenStudy (anonymous):

u there?

OpenStudy (gabylovesyou):

factor 12... 1 2 3 4 6 12

OpenStudy (gabylovesyou):

2 and 6..

OpenStudy (gabylovesyou):

(x + 2)(x + 6)

OpenStudy (anonymous):

to check ur answer expand wat u have in the brackets

OpenStudy (gabylovesyou):

x * x = x x * 6 = 6 2 * x = 2 2 * 6 = 12 im wrong lol

OpenStudy (anonymous):

the point is u know how to check ur answers now :D just try a different option :)

OpenStudy (gabylovesyou):

answer is (x - 4) (x + 3)

OpenStudy (anonymous):

YES! :D ur getting the hang of it :)